For each of the following streams of​ dividends,

Dividend Stream

Copy to Clipboard +
Open in Excel +

Year

A

B

C

20082008

​$1.801.80

20092009

​$1.861.86

20102010

​$1.931.93

20112011

​$2.002.00

​$2.182.18

20122012

​$2.072.07

​$2.242.24

20132013

​$5.205.20

​$2.142.14

​$2.292.29

20142014

​$6.016.01

​$2.212.21

​$2.322.32

20152015

​$6.816.81

​$2.282.28

​$2.372.37

20162016

​$7.597.59

​$2.332.33

​$2.412.41

20172017

​$8.418.41

​$2.452.45

​$2.502.50

​, estimate the compound annual rate of growth between the earliest year for which a value is given and 2017.

The compound annual growth rate of dividend stream A is
Answer #1

To calculate the compound annual growth rate of dividend, we use future value formula.

A) Future value( FV) = Present value( PV )*((1+growth rate( R))^no. Of years (N))

Y0 = 2013 ... Y4=2017

8.41= 5.20*(1+R)^4

R=12.77%

B) Future value( FV) = Present value( PV )*((1+growth rate( R))^no. Of years (N))

Y0 = 2008 ...Y9 = 2017

2.45= 1.80*(1+R)^9

R=3.485%

C) Future value( FV) = Present value( PV )*((1+growth rate( R))^no. Of years (N))

Y0 = 2011 ...Y6 = 2017

2.5= 2.18*(1+R)^6

R=2.31%